OpenStudy (anonymous):

solve by making appropriate substitution 2x^1/2-13x^1/4-24=0

OpenStudy (anonymous):

Let a =x^1/4, then x^1/2=(x^1/4)^2 thus 2a^2-13a-24=0 and solve by factoring

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2a^2-16a+3a-24=0\]\[2a(a-8)+3(a-8)=0\]\[(a-8)(2a+3)=0\]\[a=8 \ or \ a=-3/2\]

OpenStudy (anonymous):

(2u +3)(u-8)=0 ==> u=-3/2 or u=8